OpenStudy (anonymous):

For questions 10–11, find the solutions of the system. 10. y = x² + 3x – 4 y = 2x + 2 (a) (–3, 6) and (2, –4) (b) (–3, –4) and (2, 6) (c) (–3, –4) and (–2, –2) (d) no solution 11. y = x² – 2x – 2 y = 4x + 5 (a) (–1, 1) and (–7, –23) (b) (–1, 1) and (7, 33) (c) (–1, 33) and (7, 1) (d) no solution thankyou!!!

OpenStudy (callisto):

For 10 y = x² + 3x – 4 -(1) y = 2x + 2 -(2) (1) - (2) 0 = x² + 3x – 4 -(2x+2) 0 = x² +x -6 Can you solve for x and y from here?

OpenStudy (anonymous):

not really

OpenStudy (callisto):

Alright , do you understand so far?

OpenStudy (anonymous):

i think

OpenStudy (callisto):

then i'll continue.. 0 = x² +x -6 0 =(x+3)(x-2) x = -3 or x=2 Put x=-3 into y = 2x + 2 y = 2(-3) + 2 = -2 Put x=2 into y = 2x + 2 y = 2(2) + 2 =6 Can you choose the answer for 10 now?

OpenStudy (callisto):

Sorry.. for the middle part, it should be like this Put x=-3 into y = 2x + 2 y = 2(-3) + 2 = -4

OpenStudy (callisto):

For 11, it's similar to 10 y = x² – 2x – 2 -(1) y = 4x + 5 -(2) (1) -(2) 0 = x² – 2x – 2 -(4x + 5) 0 = x² - 6x -7 Can you do it from here then?

OpenStudy (anonymous):

A plot is attached.